# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 2.1 Composing word vectors
There are many other composition functions that can be used to derive sentence embeddings. In the early work of [27] on phrase composition, the authors also found that multiplicative and dilation models perform well. However, there are caveats: element-wise multiplication of sparse representations results in loss of information (this effect is especially strong for more than 2 words), while dilation, just like weighted average, requires additional parameters. It should also be mentioned that max pooling is quite a popular operation to aggregate the output of the LSMT encoder, as shown in [109, 12], and also by the authors of [86], who found it to work best for all BERT layers for the STS benchmark task. According to the authors of [110], the max pooling is also competitive with boundary-based methods (such as those that use embeddings of words in the boundaries) for the representation of the text span.
@@
The input text sequence can be combined in a hierarchical fashion. The authors of [111] used a recursive autoencoder to contract two child words to a single parent one repeatedly, to arrive at a final single common parent representation of the whole sequence. Later, the authors of [112] extended this combination to specifically follow the structure of the parse tree. Such a recursive model allowed them to compose word vectors in a bottom-up approach up to a sentence level and outperform the simple averaging method. However, a drawback is that such a composition is limited to only sentences. Other authors of [41] used similar ideas for combinations, but did not learn any composition function. Instead, given a specific window length, the vectors of words in these windows are averaged, and the resulting embeddings are then max-pooled. Such a setup allows one to better preserve the spatial information.
@@
The common average operation can be generalized as a power mean, as proposed for information retrieval in [113]. The authors of [114] came up with such an idea for combining word vectors. They showed that using the document vector as a concatenation of the power mean with different power values, particularly $p=\pm\infty$ (maximum and minimum), $p=1$ (regular arithmetic mean), and other $p>0$ substantially improves the representation of the document. The other authors of [115] also proposed a long vector representation. They employ self-attention to replace the max (or average) pooling from either RNN hidden states or convolved n-grams resulting in the sentence embedding of the matrix form. However, as the authors of [116] criticize, concatenating representations is problematic. According to them, bigger embeddings will always increase the performance of the linear model on top of them. Therefore, embeddings of the same size should always be compared. In light of this, concatenation is an undesirable operation.
@@ 2.2 Reshaping representation spaces
Word and sentence vectors obtained by older word2vec [1], GloVe [17] and current state-of-the-art transformer techniques are not ideal. Therefore, additional processing techniques are applied to eliminate side effects and improve the performance of downstream tasks. In this subsection, we will describe anisotropy, a phenomenon believed to be responsible for this, the reasons for its appearance, and the techniques that attempt to improve the representation space.
@@ 2.2.1 Isotropy
The anisotropic properties of independent word embeddings were first observed by the authors of [118]. They analyzed common embedding models and found that, in all cases, word vectors share a nonzero common vector; therefore, they are not zero-centered. The authors also found that the variance ratios of the first few components derived by PCA decay nearly exponentially. Furthermore, the variances explained by the leading components “encode the frequency of the word to a significant degree”. Therefore, they proposed eliminating the common mean vector and then removing the top principal components, computed on representations from the entire vocabulary. A similar procedure was also shown in the earlier work of [60] to be a very good baseline. Here, words were first combined by a frequency-weighted average, and then just the first principal component (yet dataset-specific, as computed on the entire dataset) was removed. As shown by the all-but-the-top (ABTT) method of [118], such a simple procedure, eliminating the common mean vector and a few top dominating directions from the word vectors, greatly improves both the isotropy and the downstream task performance (for sentence tasks, a simple average of preprocessed word vectors is used). As found by [119], it is already known that the isotropy of the target embedding is critical for the alignment of embeddings, which is important for areas such as domain adaptation, word embedding evaluation, and machine translation. However, the work [118] became the main stimulus for the research direction that attempts to improve the isotropy of word embeddings.
Word-representation anisotropy is also observed in the latest contextualized models. The authors of [120] first observed this effect on the transformer machine translation model [2], while the author of [98] found that this is the case for all ELMo [106], BERT [5], and GPT-2 [121] layers. All representations for randomly sampled words from these contextualized anisotropic spaces were observed to be highly similar, far from zero average cosine similarity. Furthermore, the embeddings of any two words are positively correlated. Finally, it was observed that all word vectors in the representation space tend to occupy a narrow cone.
@@
The authors of [122] were among the first to observe that word embeddings encode a surprising degree of information about word frequencies. In [60] it was shown that simply representing sentences as a weighted average of words by their frequencies is a very competitive method. In particular, the authors computed the weighted average of the word vectors in the sentence (the weight of a word w is a/(a + p(w)) with a being a parameter and p(w) the frequency) and then removed the projections of the average vectors on their first singular vector (“common component removal”). Later, the authors of [118] suggested to remove more than one component. They found that the top PCA directions encode word frequency information; therefore, the method was proposed to null these directions. As this bias is very strong in representation space, it is inevitably related to the anisotropy phenomena.
@@
Similar encoding of frequency information is also evident in contextualized embeddings of modern transformer models. The authors of [123] investigated BERT embeddings and found that high-frequency words are all close to the origin and densely concentrated, while low-frequency words are far away and scattered sparsely. The work of [124] highlighted that anisotropy is the most pronounced in rare words. The authors of [125] also showed that the same applies to multilingual BERT embeddings: they have a biased structure towards word frequency.
@@
This bias to word frequency reveals a weakness. The authors of [126] constructed a data set using semantic relations extracted from WordNet [127] to test the semantic properties of words. Using that, they showed that the ability of BERT to understand words depends highly on their frequency, and therefore rare words are neglected. In other work of [63] it was depicted that removing the 34 most frequent tokens before averaging can greatly boost the performance on semantic textual similarity tasks. The authors of [124] conducted a simple experiment on the CNN News corpus with popular WordPiece tokenization and found that 30% of the corpus can be accounted for using 13 most frequent tokens, while to cover at least 98% of the corpus, nearly 15 000 tokens are needed. Therefore, rare words are the constituent part of modern NLP pipelines and, unfortunately, induce deficiencies.
@@
The authors of [120] proposed an explanation for generation models that related observations of the narrow cone form in the representation space and the frequencies of words. They argue that the main culprit is the way models are trained on the language modeling task. During the training process the ground-truth word embedding will be pushed towards the direction of the hidden state in order to get a larger likelihood. Meanwhile, the embeddings of most words in the vocabulary (with non- and rarely-appearing frequencies) will be pushed towards similar directions negatively correlated with most hidden states and thus end up grouped together in the local region of the embedding space. The following work of [124] complemented the given explanation by common enemies effect – the effect of the target words producing gradients of the same direction for all of the non-target words at each step of training with cross-entropy loss, and rare tokens are the most affected by it. The authors also found that the embeddings learned by GPT-2, BERT, and RoBERTa do not degenerate into a narrow cone (only appear as a cone when projected to a lower-dimensional space), but instead drift in one shared direction. Therefore, common enemies effect is argued to be the main culprit behind the anisotropy of the representation space.
@@
There is evidence from multiple authors that contextualized representations of transformer models (in particular, the BERT family) contain undesirable outliers (interestingly, abnormal dimensions were also previously observed for GloVe vectors in [128]). These are certain positions in word vectors with unusually high values. The authors of [129] called it outlier dimensions; in [130] it is referred to as outlier neurons, while the authors of [102] call it rogue dimensions. In all of these works, it is agreed that it is a significant contributor to the anisotropy of the representation space.
There is a debate about what causes outliers. The authors of [129] found that outliers are essential for good downstream performance. They regard it as a distinct model property that emerges during training and even encourage one to consider it during weight initialization. In contrast, in [102] it was found that the behavior of the model is not driven by outliers; rather, only a small subset of linguistic abilities is handled here (such as positional information; its relationship with outliers was revealed in [130]). Even more strangely, the authors of [125] found that the embedding space of the multilingual BERT model does not have outliers as the English BERT does, but is still anisotropic.
There are several options to account for outliers. The most obvious solution is to remove them. The authors of [130] showed that this brings improvements for tasks that directly use the geometry of the embeddings, such as semantic textual similarity, while in [129] it was shown that disabled outliers significantly degrade both Masked Language Modeling (MLM) loss and downstream task performance. The authors of [102] advocate that rogue dimensions be taken into account through the standardization transformation. Both options aim to reduce the dominating effect of the highest-value dimensions during the similarity measure, such as cosine similarity or Pearson correlation. On the other hand, one can argue that the similarity measure that we use, not the representation space, needs improvements.
@@
There are many works that try to improve the isotropy of representation spaces, but is that truly the right way to seek performance improvements? Some recent publications, such as [124], expressed doubts about the role of isotropy in model performance, while the authors of [131] observed even a negative relation between isotropy calibration and downstream performance. The works of [124, 117] ask whether there is truly a “narrow cone” in the embedding space. Other authors of [15] argue that the main problems are various biases and are not related to isotropy. Furthermore, the authors of [117] criticized multiple popular measures of isotropy and warned that they are misleading and inaccurate.
The authors of [132] analyzed the contextual embedding space and found isolated clusters and low-dimensional manifolds. They concluded that although the embeddings are globally anisotropic, there exists local isotropy. The authors of [133] continued this observation and applied the known preprocessing techniques of [118] not globally but locally, in clustered regions of representations. They observed that sense-level information is shadowed by structural, syntactic, and tense biases (for verbs) which their method of dominant direction removal helps to reduce. In other work [134] it was shown that during the fine-tuning phase the anisotropy becomes even worse, while, in contrast, the performance of STS increases. They found that removing the top dominant directions for fine-tuned representations becomes detrimental as the most essential information moves to reside there. These findings indicate that the representation space of transformer models is quite complicated and that a simple fixation on isotropy can destroy its native features. These also include biases, the usefulness of which may not be obvious to us, but models have found them helpful during the pre-training. In the end, brute-forcing the plain isotropy may be harmful to some intricate semantic details, which we seek to capture.
@@ 2.2.2 Post-processing methods
Many post-processing methods have been proposed to improve the embeddings. Whether they improve the isotropy of the representation space or help surface relevant information [135], multiple authors show the benefits of using them. However, these methods tend to generally favor unsupervised tasks, such as semantic textual similarity, with limited improvements for downstream tasks. If enough data on the target task are available, supervised training will inevitably be a better choice, as the transformation relevant to the task is learned. However, if training samples are obscure, post-processing may be the only option.
@@
The simplest is the centering operation. It is especially important for the STS task, where cosine similarities are involved. The authors of [124] found that it restores a nearly perfectly isotropic distribution. Subtracting the mean is also the first step in several other post-processing methods. In some sense, such operations as centering can be viewed as fine-tuning to the target domain, as all target embeddings participate in providing the mean coordinates.
@@
We already mentioned some operations related to isotropy enhancement, such as zeroing target outliers or ABTT, in the previous section. In the following, we will try to mention the remaining important ones that were left untouched.
z-score
@@
Also called standardization, it describes the position of a point in terms of its distance from the mean when measured in units of standard deviation (the distribution is that of the target samples). In other words, it transforms the distribution of embeddings to have a zero mean and a unit standard deviation in all dimensions. The works of [102, 103] recommend that you consider it as a post-processing step. The authors of [116] advise using normalization, such as z-score, in supervised settings as a binary hyperparameter, because it can lead to rank changes.
All-but-the-top
@@
Many works are based on the influential ABTT algorithm of [118]. It has a hyperparameter, which tells the number of dominant directions to remove. Principle components are either completely removed or left intact. As a result, the main weakness is the decision on the right number of top components to remove, which is prone to either cause information loss or eliminate an insufficient amount of noise. The authors of [136] proposed post-processing through variance normalization (PVN) to normalize the variance of leading principal components to the same level instead of total component disposal. The authors of [137] performed the removal in a softer way by employing matrix conceptors [138] in an unsupervised way. A similar technique that employs conceptors can also be used to target specific components, such as those that incorporate social biases and diminish them, as in [139]. Other authors of [140] proposed learning weights for each dominant direction removal. All these approaches try to preserve useful information that resides in the top principal components while narrowing the target noise.
@@
Authors of [141] combined all-but-the-top and dimensionality reduction. They found that the best pipeline is to perform ABTT twice while performing dimensionality reduction in between. In another line of work, the authors of [133, 125], instead of global post-processing, remove local dominant directions in separate clusters in the representation space. In this way, the structure of the representation space is taken into account. Similarly, the authors of [65], in light of the predecessor method of ABTT, the work of [60], proposed to eliminate the sentence-dependent principal component, where they rerank the top principal vectors based on correlation with each sentence. This individual removal of dominant directions was shown to improve performance on the STSB task.
@@
The authors of [13] proposed the whitening approach to alleviate the anisotropy problem of sentence embedding. The whitening operation involves centering embeddings at the origin and making different dimensions have a unit variance and be uncorrelated among themselves, turning their covariance matrices into the identity matrix. It was also found to be useful in earlier work for the alignment of bilingual embeddings [142]. As a side effect, the authors of [13] showed that whitening can also be used with dimension reduction operation. A concurrent work of [14] also saw a whitening algorithm to improve performance on STS tasks. In that work, the authors first combined the first and last layers of BERT embeddings and then normalized them with whitening. Despite the initial success, subsequent work criticized the whitening operation. The authors of [143] showed that whitened representations greatly improve uniformity, but also suffer degeneration in the alignment property (for alignment and uniformity properties see [144]). To make matters worse, the authors of [145] called whitening a “trick” that only helps with similarity tasks (by partially overfitting) and harms downstream task performance.

### v2

@@ 2.1.9 Other Means of Composition
Many other composition functions can be used to derive sentence embeddings. In an early work [27] on phrase composition, the authors also found that multiplicative and dilation models perform well. However, there are caveats; element-wise multiplication of sparse representations results in loss of information (this effect is especially strong for more than two words), while dilation, just like weighted average, requires additional parameters. It should also be mentioned that max pooling is quite a popular operation to aggregate the output of the LSMT encoder, as shown in [109, 12], and also by the authors of [86], who found it to work best for all BERT layers for the STS benchmark task. According to the authors of [110], max pooling is also competitive with boundary-based methods (such as those that use embeddings of words in the boundaries) for the representation of the text span.
@@
The input text sequence can be combined hierarchically. The authors of [111] used a recursive autoencoder to contract two child words to a single parent one repeatedly, to arrive at a final single common parent representation of the whole sequence. Later, the authors of [112] extended this combination to specifically follow the structure of the parse tree. Such a recursive model allowed them to compose word vectors in a bottom-up approach up to a sentence level and outperform the simple averaging method. However, a drawback is that such a composition is limited to only sentences. Other authors of [41] used similar ideas for combinations, but did not learn any composition function. Instead, given a specific window length, the vectors of words in these windows are averaged, and the resulting embeddings are then max-pooled. Such a setup better preserves the spatial information.
@@
The common average operation can be generalized as a power mean, as proposed for information retrieval in [113]. The authors of [114] introduced such an idea for combining word vectors. They showed that using the document vector as a concatenation of the power mean with different power values, particularly $p=\pm\infty$ (maximum and minimum), $p=1$ (regular arithmetic mean), and other $p>0$, substantially improves the representation of the document. Other authors of [115] also proposed a long vector representation. They employ self-attention to replace the max (or average) pooling from either RNN hidden states or convolved n-grams, resulting in the sentence embedding of the matrix form. However, as the authors of [116] criticize, concatenating representations is problematic. According to them, bigger embeddings will always increase the performance of the linear model on top of them. Therefore, embeddings of the same size should always be compared. In light of this, concatenation is an undesirable operation.
@@ 2.2 Reshaping Representation Spaces
Word and sentence vectors obtained by older Word2Vec [1], GloVe [17], and current state-of-the-art transformer techniques are not ideal. Therefore, additional processing techniques are applied to eliminate side effects and improve the performance of downstream tasks. In this subsection, we will describe anisotropy, a phenomenon believed to be responsible for this, the reasons for its appearance, and the techniques that attempt to improve the representation space.
@@ 2.2.1 Isotropy
The anisotropic properties of independent word embeddings were first observed by the authors of [118]. They analyzed common embedding models and found that, in all cases, word vectors share a nonzero common vector; therefore, they are not zero-centered. The authors also found that the variance ratios of the first few components derived by PCA decay nearly exponentially. Furthermore, the variances explained by the leading components “encode the frequency of the word to a significant degree”. Therefore, they proposed eliminating the common mean vector and then removing the top principal components, computed on representations from the entire vocabulary. A similar procedure was also shown in an earlier work [60] to be a very good baseline. Here, words were combined by a frequency-weighted average, and then just the first principal component (yet dataset-specific, as computed on the entire dataset) was removed. As shown by the All-But-The-Top (ABTT) method of [118], such a simple procedure, eliminating the common mean vector and a few top dominating directions from the word vectors, greatly improves both the isotropy and the downstream task performance (for sentence tasks, a simple average of preprocessed word vectors is used). As found by [119], it is already known that the isotropy of the target embedding is critical for the alignment of embeddings, which is important for areas such as domain adaptation, word embedding evaluation, and machine translation. However, the work [118] became the main stimulus for the research direction that attempts to improve the isotropy of word embeddings.
Word representation anisotropy is also observed in the latest contextualized models. The authors of [120] first observed this effect on the transformer machine translation model [2], while the author of [98] found that this is the case for all ELMo [106], BERT [5], and GPT-2 [121] layers. All representations for randomly sampled words from these contextualized anisotropic spaces were observed to be highly similar, far from zero average cosine similarity. Furthermore, the embeddings of any two words are positively correlated. Finally, it was observed that all word vectors in the representation space tend to occupy a narrow cone.
@@
The authors of [122] were among the first to observe that word embeddings encode a surprising degree of information about word frequencies. In [60], it was shown that simply representing sentences as a weighted average of words by their frequencies is a very competitive method. In particular, the authors computed the weighted average of the word vectors in the sentence (the weight of a word $w$ is $a/(a+p(w))$, with $a$ being a parameter and $p(w)$ the frequency) and then removed the projections of the average vectors on their first singular vector (“common component removal”). Later, the authors of [118] suggested removing more than one component. They found that the top PCA directions encode word frequency information; therefore, the method was proposed to null these directions. As this bias is very strong in representation space, it is inevitably related to the anisotropy phenomena.
@@
Similar encoding of frequency information is also evident in contextualized embeddings of modern transformer models. The authors of [123] investigated BERT embeddings and found that high-frequency words are all close to the origin and densely concentrated, while low-frequency words are far away and scattered sparsely. The work in [124] highlighted that anisotropy is the most pronounced in rare words. The authors of [125] also showed that the same applies to multilingual BERT embeddings; they have a biased structure towards word frequency.
@@
This bias to word frequency reveals a weakness. The authors of [126] constructed a dataset using semantic relations extracted from WordNet [127] to test the semantic properties of words. Using that, they showed that the ability of BERT to understand words depends highly on their frequency, and therefore rare words are neglected. It was depicted in [63] that removing the 34 most frequent tokens before averaging can greatly boost the performance on semantic textual similarity tasks. The authors of [124] conducted a simple experiment on the CNN News corpus with popular WordPiece tokenization and found that 30% of the corpus can be accounted for using the 13 most frequent tokens, while to cover at least 98% of the corpus, nearly 15,000 tokens are needed. Therefore, rare words are a constituent part of modern NLP pipelines and, unfortunately, induce deficiencies.
@@
The authors of [120] proposed an explanation for generation models that related observations of the narrow cone form in the representation space and the frequencies of words. They argue that the main culprit is the way models are trained on the language modeling task. During the training process, the ground-truth word embedding will be pushed toward the direction of the hidden state to obtain a larger likelihood. Meanwhile, the embeddings of most words in the vocabulary (with non-appearing and rarely appearing frequencies) will be pushed towards similar directions negatively correlated with most hidden states and thus end up grouped together in the local region of the embedding space. The following work [124] complemented the given explanation with the “common enemies effect”—the effect of the target words producing gradients of the same direction for all of the non-target words at each step of training with cross-entropy loss, and rare tokens are the most affected by it. The authors also found that the embeddings learned by GPT-2, BERT, and RoBERTa do not degenerate into a narrow cone (they only appear as a cone when projected to a lower-dimensional space), but instead drift in one shared direction. Therefore, the “common enemies effect” is argued to be the main culprit behind the anisotropy of the representation space.
@@
There is evidence from multiple authors that contextualized representations of transformer models (in particular, the BERT family) contain undesirable outliers (interestingly, abnormal dimensions were also previously observed for GloVe vectors in [128]). These are certain positions in word vectors with unusually high values. The authors of [129] called it outlier dimensions; in [130], it is referred to as outlier neurons, while the authors of [102] call it rogue dimensions. In all of these works, it is agreed that it is a significant contributor to the anisotropy of the representation space.
There is debate about what causes outliers. The authors of [129] found that outliers are essential for good downstream performance. They regard it as a distinct model property that emerges during training and even encourage one to consider it during weight initialization. In contrast, in [102], it was found that the behavior of the model is not driven by outliers; rather, only a small subset of linguistic abilities is handled there (such as positional information; its relationship with outliers was revealed in [130]). Even more strangely, the authors of [125] found that the embedding space of the multilingual BERT model does not have outliers as the English BERT does, but is still anisotropic.
There are several options to account for outliers. The most obvious solution is to remove them. The authors of [130] showed that this brings improvements for tasks that directly use the geometry of the embeddings, such as semantic textual similarity; however, ref. [129] showed that disabled outliers significantly degrade both Masked Language Modeling (MLM) loss and downstream task performance. The authors of [102] suggest that rogue dimensions be accounted for through the standardization transformation. Both options aim to reduce the dominating effect of the highest-value dimensions during the similarity measure, such as cosine similarity or Pearson correlation. On the other hand, one can argue that the similarity measure we use, not the representation space, needs improvements.
@@
Many works try to improve the isotropy of representation spaces, but is that truly the right way to seek performance improvements? Some recent publications, such as [124], expressed doubts about the role of isotropy in model performance, and the authors of [131] even observed a negative relation between isotropy calibration and downstream performance. The works in [124, 117] ask whether there is truly a “narrow cone” in the embedding space. Other authors of [15] argue that the main problems are various biases unrelated to isotropy. Furthermore, the authors of [117] criticized multiple popular measures of isotropy and warned that they are misleading and inaccurate.
The authors of [132] analyzed the contextual embedding space and found isolated clusters and low-dimensional manifolds. They concluded that although the embeddings are globally anisotropic, local isotropy exists. The authors of [133] continued this observation and applied the known preprocessing techniques of [118], not globally but locally, in clustered regions of representations. They observed that sense-level information is shadowed by structural, syntactic, and tense biases (for verbs), which their method of dominant direction removal helps to reduce. Another work [134] showed that during the fine-tuning phase, the anisotropy becomes even worse, while, in contrast, the performance of STS increases. They found that removing the top dominant directions for fine-tuned representations becomes detrimental, as the most essential information resides there. These findings indicate that the representation space of transformer models is quite complicated and that a simple fixation on isotropy can destroy its native features. These also include biases, the usefulness of which may not be obvious to us, but models have found them helpful during the pre-training. In the end, brute-forcing the plain isotropy may harm some intricate semantic details, which we seek to capture.
@@ 2.2.2 Post-Processing Methods
Many post-processing methods have been proposed to improve the embeddings. Whether they improve the isotropy of the representation space or help surface-relevant information [135], multiple authors show the benefits of using them. However, these methods generally favor unsupervised tasks, such as semantic textual similarity, with limited improvements for downstream tasks. If enough data on the target task are available, supervised training will inevitably be a better choice, as the transformation relevant to the task is learned. However, post-processing may be the only option if the training samples are obscure.
@@
The simplest is the centering operation. It is especially important for the STS task that involves cosine similarities. The authors of [124] found that it restores a nearly perfectly isotropic distribution. Subtracting the mean is also the first step in several other post-processing methods. In some sense, such operations as centering can be viewed as fine-tuning to the target domain, as all target embeddings participate in providing the mean coordinates.
@@
We already mentioned some operations related to isotropy enhancement in the previous section, such as zeroing target outliers or ABTT. In the following, we will mention the remaining important ones.
Z-Score Normalization
@@
The Z-score describes the position of a point in terms of its distance from the mean when measured in units of standard deviation (in the distribution of the target samples). Z-score normalization, also called standardization, transforms the distribution of embeddings to have a zero mean and a unit standard deviation in all dimensions. The works in [102, 103] recommend that you consider it as a post-processing step. The authors of [116] advise using normalization, such as the z-score, in supervised settings as a binary hyperparameter, because it can lead to rank changes.
All-But-The-Top (ABTT)
@@
Many works are based on the influential ABTT algorithm [118]. It has a hyperparameter, which tells the number of dominant directions to remove. Principle components are either completely removed or left intact. As a result, the main weakness is finding the right number of top components to remove, which can either cause information loss or eliminate an insufficient amount of noise. The authors of [136] proposed post-processing through variance normalization (PVN) to normalize the variance of leading principal components to the same level instead of total component disposal. The authors of [137] performed the removal in a softer way by employing matrix conceptors [138] in an unsupervised way. A similar technique that employs conceptors can also target specific components, such as those that incorporate social biases, and diminish them, as in [139]. Other authors of [140] proposed learning weights for each dominant direction removal. All these approaches try to preserve useful information residing in the top principal components while narrowing the target noise.
@@
Authors of [141] combined ABTT and dimensionality reduction. They found that the best pipeline is to perform ABTT twice while performing dimensionality reduction in between. In another line of work, the authors of [133, 125], instead of global post-processing, remove local dominant directions in separate clusters in the representation space. In this way, the structure of the representation space is accounted for. Similarly, the authors of [65], in light of the predecessor method of ABTT [60], proposed to eliminate the sentence-dependent principal component, where they rerank the top principal vectors based on correlation with each sentence. This individual removal of dominant directions was shown to improve performance on the STSB task.
@@
The authors of [13] proposed the whitening approach to alleviate the anisotropy problem of sentence embedding. The whitening operation involves centering embeddings at the origin and making different dimensions have a unit variance and be uncorrelated, turning their covariance matrix into the identity matrix. It was also found to be useful in earlier work for the alignment of bilingual embeddings [142]. As a side effect, the authors of [13] showed that whitening can also be used with a dimension reduction operation. A concurrent work [14] also used a whitening algorithm to improve performance on STS tasks. In that work, the authors combined the first and last layers of BERT embeddings and then normalized them with whitening. Despite the initial success, subsequent work criticized the whitening operation. The authors of [143] showed that whitened representations greatly improve uniformity, but also suffer degeneration in the alignment property (for alignment and uniformity properties, see [144]). To make matters worse, the authors of [145] called whitening a “trick” that only helps with similarity tasks (by partially overfitting) and harms downstream task performance.